# High-resolution radio imaging of TGSS J1530++1049, a radio galaxy in a dense environment at $z=4$

### v1

@@ Abstract
Context. High-redshift radio galaxies can provide important insights into the structure formation and galaxy evolution at earlier cosmological epochs. TGSS J1530++1049 was selected as a candidate high-redshift radio galaxy, based on its very steep radio spectrum. Subsequent observations with the James Webb Space Telescope (JWST) presented in a companion paper (Saxena et al., submitted) have shown that it is located at a redshift $z=4.0$. The JWST data furthermore showed that the radio source is part of one of the densest structures of galaxies and ionized gas known at these redshifts. The complex system qualitatively resembles a massive (cluster) galaxy forming early through a rapid succession of mergers.
@@ 1 Introduction
Historically, it was proposed that compact radio sources with ultra-steep radio spectra could be good candidates for high-redshift radio galaxies (e.g., Blumenthal_Miley; Rottgering1994; Chambers_1996). Various authors use different definitions for the ultra-steep-spectrum (USS) sources (see, e.g., Brude+1995Ap&SS, and, for a later summary, Coppejans2017), but more recently Saxena_sample collected a sample of USS sources with radio spectral indices $\alpha<-1.3$ ($\alpha$ is defined as $S\propto\nu^{\alpha}$, where $S$ is the flux density and $\nu$ the observing frequency) measured between $150$ MHz and $1.4$ GHz. To enable further efficient selection of potential high-redshift sources, Saxena_sample focused on those which remained undetected in optical and infrared surveys.
TGSS J1530++1049 (hereafter J1530++1049) was one of the brightest objects in their sample. In a follow-up radio observation of the sample, conducted with the Karl G. Jansky Very Large Array (VLA) at $1.4$ GHz in its most extended A configuration, J1530++1049 was detected as a source well-described by a single Gaussian brightness distribution with a deconvolved size of $\sim 0.6\arcsec$, and a flux density of $(7.5\pm 0.1)$ mJy. Optical spectroscopic observation at the radio position revealed a single emission line that was erroneously identified as a Lyman-$\alpha$ line at $z=5.72$ (disc). However, recent James Webb Space Telescope (JWST) imaging and spectroscopic observations of the object presented in a companion paper (Saxena et al. submitted) have revealed that the radio emission positionally coincides with a highly complex system consisting of several massive galaxies, diffuse emission and ionized gas knots at $z=4.0$. The picture that has emerged is that of a dense core of interacting galaxies including the host galaxy of the radio source, perhaps signaling the formation of a very massive (central cluster) galaxy.
We observed J1530++1049 with the technique of very long baseline interferometry (VLBI) using the European VLBI Network (EVN) and the enhanced Multi-Element Remotely Linked Interferometer Network (e-MERLIN) to resolve its radio emission. Preliminary results of the EVN observation were presented in rnaas, which used the redshift estimate available at that time. In Sect. 2, we describe the observations and data reduction. Results are presented in Sect. 3, while our findings are discussed in Sect. 4. The paper is concluded by a summary in Sect. 5.
In the following, we assume a $\Lambda$ Cold Dark Matter cosmological model with a Hubble constant of $H_{0}=67.7\,\,\mathrm{km\,s}^{-1}\mathrm{\,Mpc}^{-1}$, matter density parameter of $\Omega_{\mathrm{m}}=0.31$, and dark energy density parameter of $\Omega_{\Lambda}=0.69$. At $z=4.0$, the angular size of $1$ milliarcsecond (mas) corresponds to $7.1$ pc linear size, and the luminosity distance of the source is $D_{\mathrm{L}}=36676.4$ Mpc (Cosmocalc).
@@ 2.1 EVN observation
The EVN observation of J1530++1049 was conducted on 2018 September 19 at $1.7$ GHz frequency (project code: RSG11) in e-VLBI mode (evlbi). The observing bandwidth was divided into $8$ intermediate frequency channels (IF) with $16$ MHz bandwidth each. The following antennas provided data: a single antenna of the Westerbork Synthesis Radio Telescope (the Netherlands), Effelsberg (Germany), Medicina (Italy), Onsala (Sweden), Tianma (China), Toruń (Poland), Hartebeesthoek (South Africa), and Sardinia (Italy). Sardinia observed in only four IFs.
The observation was conducted in phase-referencing mode (p-ref). The phase-reference calibrator, ICRF J152502.9++110744 (hereafter J1525++1107), and the target source were observed alternately, in cycles with $1.5$ min time spent on the calibrator and $4.5$ min on the target source. The total on-target integration time was $57.6$ min.
The data were reduced with the National Radio Astronomy Observatory (NRAO) Astronomical Image Processing System (AIPS, aips) following the standard procedure (e.g., VLBI_reduc). A priori amplitude calibration was performed using system temperature measurements at the antenna sites. Ionospheric corrections were conducted based on the total electron content maps from global navigation satellite systems data, and parallactic angle correction was also done. Then fringe fitting was performed on the data of the phase-reference calibrator.
The resulting calibrator source data were imaged in the Difmap software (difmap) via the hybrid mapping method, which involves subsequent usage of the CLEAN algorithm (clean) and phase self-calibration iterations. As a final step of imaging, we performed amplitude self-calibration and obtained overall gain correction factors for the anntennas. These gain corrections were then applied to the data within AIPS. Additionally, the resulting image of the phase-reference calibrator was read back into AIPS to be used for improving the fringe fit. The fringe-fit solutions of the phase-reference calibrator were applied to the target source, which was than imaged in Difmap. Due to the weakness of the main target object, no self-calibration was attempted.
@@
The e-MERLIN observations (project code: CY8205) took place on 2019 January 29 and February 22 at $5$ GHz, and on 2019 March 7 and 8 at $1.5$ GHz. At $1.5$ GHz, $8$ IFs were used, each with a bandwidth of $64$ MHz and divided into $128$ spectral channels. However, no signal was recorded in IF77. At $5$ GHz, $4$ IFs were used, each with a bandwidth of $128$ MHz and divided into $128$ channels. At both frequencies, the Jodrell Bank Mk2, Pickmere, Darnhall, Knockin, Defford, and Cambridge antennas participated in the observations. The Lovell Telescope at Jodrell Bank also joined at $5$ GHz on the second observing day.
@@
The raw visibility data were pre-processed and calibrated with the e-MERLIN pipeline111version 0.9 using the Common Astronomy Software Applications (CASA, casa_old; casa)222version 5.4. Imaging was also performed in CASA. We exported the channel-averaged dataset and additionally imaged it in Difmap. The resulting maps agree with each other. Due to the weakness of the source, no self-calibration was performed.
@@ 3.1 EVN observation
Two compact features (labeled as A and B) were detected with the EVN at $1.7$ GHz with a projected separation of $\sim 400$ mas (Fig. 1), corresponding to $\sim 2.8$ kpc at the redshift of the source.
The coordinates of the brighter, northern feature (A) are right ascension $\alpha=15^{\mathrm{h}}30^{\mathrm{m}}49\aas@@fstack{s}89029\pm 0\aas@@fstack{s}00013$, and declination $\delta=10\degr 49\arcmin 31\aas@@fstack{\prime\prime}1754\pm 0\aas@@fstack{\prime\prime}002$. When calculating the positional error, we took into account the uncertainty of the coordinates of the phase-reference calibrator source ($0.11$ mas in right ascension and $0.18$ mas in declination direction [^3] According to the latest Radio Fundamental Catalog (rfc) rfc2025b maintained by L. Petrov at https://astrogeo.org/sol/rfc/rfc_2025b/), and the error arising from the phase-referencing technique which depends on the target and calibrator source separation, in our case $1\aas@@fstack{\circ}45$. multiview estimate this error to be $\sim 2$ mas. pradel_charlot obtained a close value of $1.5$ mas via simulated VLBI observations. We used the larger, more conservative value. It should also be noted that the coordinates of the phase-reference calibrator source were measured at $8.4$ GHz, while the EVN observation was conducted at $1.7$ GHz. Frequency-dependent core shift (coreshift) could cause a difference between the International Celestial Reference Frame (ICRF, icrf3) position determined at $8.4$ GHz and the $1.7$-GHz position of the calibrator.
@@
To describe the emission quantitatively, we fitted the visibilities with circular Gaussian brightness distribution model components. Since self-calibration cannot be performed, the obtained flux densities are expected to be lower than the real value due to coherence loss (coherence-loss). Previous works (e.g. mosoni_coherenceloss; gabanyi_coherenceloss) estimated the coherence loss to be $\sim 25$ %. The parameters of the two fitted components are given in Table 1, where the flux densities have been increased to account for the coherence loss. The errors were calculated according to the formulae given by error_est. For the flux density errors, another $10$ % was added in quadrature to account for the absolute amplitude calibration inaccuracies.
The brightness temperature of the features can be calculated as (e.g., tb_veres)
@@
where $S$ is the flux density of the components measured in Jy, $\nu$ the observing frequency in GHz, and $\theta$ the full width at half-maximum (FWHM) size in mas. The values obtained are $T_{\mathrm{B}}^{\mathrm{A}}=(4.0\pm 0.8)\cdot 10^{7}$ K and $T_{\mathrm{B}}^{\mathrm{B}}=(8.2\pm 3.2)\cdot 10^{7}$ K for components A and B, respectively. Both values well exceed the upper limit given for star-forming galaxies by condon_radio, $\sim 10^{5}$ K, indicating AGN-related radio emission.
@@
Figure 3: 5-GHz e-MERLIN image of J1530+1049. The peak intensity is $0.17\mathrm{\,mJy\,beam}^{-1}$. The lowest contours are drawn at $\pm 0.046\mathrm{\,mJy\,beam}^{-1}$ corresponding to $3.5\sigma$ image noise level. Further positive contour levels increase by a factor of $\sqrt{2}$. The restoring beam is shown in the upper left corner, its FWHM size is $121\mathrm{\,mas}\times 36$ mas and its major axis is oriented at a position angle of $23\degr$. The pink crosses indicate the positions of the EVN-detected components.
With e-MERLIN, two features (designated as N and S) could be detected at $1.5$ GHz with a separation of $\sim 780$ mas, as shown in Fig. 2. At $5$ GHz, the brighter northern feature (N) is clearly detected with e-MERLIN. At the approximate distance of feature S, at $\sim 700$ mas from N, and in the same position angle as seen in Fig. 2, we recover an emission spot at $\sim 5\sigma$ image noise level at $5$ GHz (Fig. 3).
@@
The results of the model fitting are given in Table 2. The errors have been estimated using the formulae of error_est, and additional $20\%$ and $10\%$ were added quadratically to the flux density errors to account for the amplitude calibration uncertainty of e-MERLIN at $1.5$ GHz and $5$ GHz, respectively (Lband_error; Cband_error).
@@
Figure 4: The flux densities of the two Gaussian components that can be fitted to the e-MERLIN visibilities of J1530++1049. Circles and squares are for the components N and S, respectively. The solid and dashed lines are power-law fits to the corresponding data points. For comparison, the sum of the flux densities of the EVN-detected components is also shown with a yellow upward triangle. Empty downward brown triangles show lower-resolution flux density measurements from various surveys (see Sect. 4.3 for details), and the brown line indicates the power-law fit to those points. The numbers show the corresponding spectral indices.
@@ 4.1 Astrometric registration of the radio components
The same phase-reference source (J1525++1107) was used in all observations. However, it has significant radio emission extended to $\sim 100$-mas scale and the position of its brightness peak can be different at different observing frequencies. To illustrate this, the e-MERLIN images of the calibrator are shown in Fig. 5, with a shift of $\sim 110$ mas between the image cutouts to facilitate their comparison. The north–south oriented double feature seen at $5$ GHz cannot be resolved at $1.5$ GHz with e-MERLIN. However, if the flux densities of the two components relative to each other change with frequency, i.e. the northern feature has higher or very similar flux density than the southern one at the lower frequency of $1.5$ GHz, then the position of the brightness peak can shift towards the northern feature which, on the other hand, is relatively fainter at $5$ GHz. This frequency-dependent change in the location of the brightness peak is further supported by higher-resolution VLBI images of the calibrator taken at $2.3$ GHz, showing an extended bright feature to the north of the mas-scale compact emitting region (the core) at a distance of $\sim(120-150)$ mas. These archival images are available at the Astrogeo website [^4] http://astrogeo.org/cgi-bin/imdb_get_source.csh?source=J1525%2B1107 maintained by L. Petrov. Thus we can conclude that the positional mismatch between the $1.5$-GHz and $5$-GHz e-MERLIN images of J1530++1049 is caused by different brightness distributions of the phase-reference calibrator source at the two frequencies used (Fig. 5). Therefore, in our target source, the brighter northern features detected by e-MERLIN at the two frequencies (Figs. 2 and 3) are co-spatial, and the brighter EVN-detected component (A, Fig. 1) is located within this emitting region. Since the distances and position angles of the target morphology are very similar in the two e-MERLIN images, and no other significant radio emission can be detected in the field at either of the frequencies, we conclude that their cross-identification is solid.

### v2

@@ Abstract
Context. High-redshift radio galaxies can provide important insights into the structure formation and galaxy evolution at earlier cosmological epochs. TGSS J1530++1049 was selected as a candidate high-redshift radio galaxy, based on its very steep radio spectrum. Subsequent observations with the James Webb Space Telescope (JWST) presented in a companion paper have shown that it is located at a redshift $z=4.0$. The JWST data furthermore showed that the radio source is part of one of the densest structures of galaxies and ionized gas known at these redshifts. The complex system qualitatively resembles a massive (cluster) galaxy forming early through a rapid succession of mergers.
@@ 1 Introduction
Historically, it was proposed that compact radio sources with ultra-steep radio spectra could be good candidates for high-redshift radio galaxies (e.g., Blumenthal and Miley 1979; Roettgering et al. 1994; Chambers et al. 1996). Various authors use different definitions for the ultra-steep-spectrum (USS) sources (see, e.g., Braude et al. (1995), and, for a later summary, Coppejans et al. (2017)), but more recently Saxena et al. (2018a) collected a sample of USS sources with radio spectral indices $\alpha<-1.3$ ($\alpha$ is defined as $S\propto\nu^{\alpha}$, where $S$ is the flux density and $\nu$ the observing frequency) measured between $150$ MHz and $1.4$ GHz. To enable further efficient selection of potential high-redshift sources, Saxena et al. (2018a) focused on those which remained undetected in optical and infrared surveys.
TGSS J1530++1049 (hereafter J1530++1049) was one of the brightest objects in their sample. In a follow-up radio observation of the sample, conducted with the Karl G. Jansky Very Large Array (VLA) at $1.4$ GHz in its most extended A configuration, J1530++1049 was detected as a source well-described by a single Gaussian brightness distribution with a deconvolved size of $\sim 0.6\arcsec$, and a flux density of $(7.5\pm 0.1)$ mJy. Optical spectroscopic observation at the radio position revealed a single emission line that was erroneously identified as a Lyman-$\alpha$ line at $z=5.72$ (Saxena et al. 2018b). However, recent James Webb Space Telescope (JWST) imaging and spectroscopic observations of the object presented in a companion paper (Saxena et al. 2026) have revealed that the radio emission positionally coincides with a highly complex system consisting of several massive galaxies, diffuse emission and ionized gas knots at $z=4.0$. The picture that has emerged is that of a dense core of interacting galaxies including the host galaxy of the radio source, perhaps signaling the formation of a very massive (central cluster) galaxy.
We observed J1530++1049 with the technique of very long baseline interferometry (VLBI) using the European VLBI Network (EVN) and the enhanced Multi-Element Remotely Linked Interferometer Network (e-MERLIN) to resolve its radio emission. Preliminary analysis of the EVN observation were presented in Gabányi et al. (2018), which used the redshift estimate available at that time. In Sect. 2, we describe the observations and data reduction. Results are presented in Sect. 3, while our findings are discussed in Sect. 4. The paper is concluded by a summary in Sect. 5.
In the following, we assume a $\Lambda$ Cold Dark Matter cosmological model with a Hubble constant of $H_{0}=67.7\,\,\mathrm{km\,s}^{-1}\mathrm{\,Mpc}^{-1}$, matter density parameter of $\Omega_{\mathrm{m}}=0.31$, and dark energy density parameter of $\Omega_{\Lambda}=0.69$. At $z=4.0$, the angular size of $1$ milliarcsecond (mas) corresponds to $7.1$ pc linear size, and the luminosity distance of the source is $D_{\mathrm{L}}=36676.4$ Mpc (Wright 2006).
@@ 2 Observations and data reduction
Figure 1: The $(u,v)$-coverages of the three radio interferometric observations of J1530++1049, from left to right, the $1.5$-GHz e-MERLIN, the $5$-GHz e-MERLIN, and the $1.7$-GHz EVN observations. We note that the units of the axes are $10^{3}\lambda$ for the e-MERLIN observations and $10^{6}\lambda$ for the EVN observation, where $\lambda$ indicates the observing wavelength.
@@ 2.1 EVN observation
The EVN observation of J1530++1049 was conducted on 2018 September 19 at $1.7$ GHz frequency (project code: RSG11) in e-VLBI mode (Szomoru 2008). The observing bandwidth was divided into $8$ intermediate frequency channels (IF) with $16$ MHz bandwidth each. The following antennas provided data: a single antenna of the Westerbork Synthesis Radio Telescope (the Netherlands), Effelsberg (Germany), Medicina (Italy), Onsala (Sweden), Tianma (China), Toruń (Poland), Hartebeesthoek (South Africa), and Sardinia (Italy). Sardinia observed in only four IFs. The $(u,v)$-coverage of the observation is shown in the rightmost panel of Fig. 1.
The observation was conducted in phase-referencing mode (Beasley and Conway 1995). The phase-reference calibrator, ICRF J152502.9++110744 (hereafter J1525++1107), and the target source were observed alternately, in cycles with $1.5$ min time spent on the calibrator and $4.5$ min on the target source. The total on-target integration time was $57.6$ min.
The data were reduced with the National Radio Astronomy Observatory (NRAO) Astronomical Image Processing System (AIPS, Greisen 1990) following the standard procedure (e.g., Diamond 1995). A priori amplitude calibration was performed using system temperature measurements at the antenna sites. Ionospheric corrections were conducted based on the total electron content maps from global navigation satellite systems data, and parallactic angle correction was also done. Then fringe fitting was performed on the data of the phase-reference calibrator.
The resulting calibrator source data were imaged in the Difmap software (Shepherd 1997) via the hybrid mapping method, which involves subsequent usage of the CLEAN algorithm (Högbom 1974) and phase self-calibration iterations. As a final step of imaging, we performed amplitude self-calibration and obtained overall gain correction factors for the antennas. These gain corrections were then applied to the data within AIPS. Additionally, the resulting image of the phase-reference calibrator was read back into AIPS to be used for improving the fringe fit. The fringe-fit solutions of the phase-reference calibrator were applied to the target source, which was than imaged in Difmap. Due to the weakness of the main target object, no self-calibration was attempted.
@@
The e-MERLIN observations (project code: CY8205) took place on 2019 January 29 and February 22 at $5$ GHz, and on 2019 March 7 and 8 at $1.5$ GHz. At $1.5$ GHz, $8$ IFs were used, each with a bandwidth of $64$ MHz and divided into $128$ spectral channels. However, no signal was recorded in IF77. At $5$ GHz, $4$ IFs were used, each with a bandwidth of $128$ MHz and divided into $128$ channels. At both frequencies, the Jodrell Bank Mk2, Pickmere, Darnhall, Knockin, Defford, and Cambridge antennas participated in the observations. The Lovell Telescope at Jodrell Bank also joined at $5$ GHz on the second observing day. The $(u,v)$-coverages of the observations are shown in the left and middle panels of Fig. 1.
@@
The raw visibility data were pre-processed and calibrated with the e-MERLIN pipeline111version 0.9 using the Common Astronomy Software Applications (CASA, McMullin et al. 2007; CASA Team et al. 2022)222version 5.4. Imaging was also performed in CASA. We exported the channel-averaged dataset and additionally imaged it in Difmap. The resulting maps agree with each other. Due to the weakness of the source, no self-calibration was performed.
@@ 3.1 EVN observation
Two mas-scale compact features (labeled as A and B) were detected with the EVN at $1.7$ GHz with a projected separation of $\sim 400$ mas (Fig. 2), corresponding to $\sim 2.8$ kpc at the redshift of the source.
The coordinates of the brighter, northern feature (A) are right ascension $\alpha=15^{\mathrm{h}}30^{\mathrm{m}}49\aas@@fstack{s}89029\pm 0\aas@@fstack{s}00013$, and declination $\delta=10\degr 49\arcmin 31\aas@@fstack{\prime\prime}1754\pm 0\aas@@fstack{\prime\prime}002$. When calculating the positional error, we took into account the uncertainty of the coordinates of the phase-reference calibrator source ($0.11$ mas in right ascension and $0.18$ mas in declination direction [^3] According to the latest Radio Fundamental Catalog (Petrov and Kovalev 2025) rfc2025b maintained by L. Petrov at https://astrogeo.org/sol/rfc/rfc_2025b/), and the error arising from the phase-referencing technique which depends on the target and calibrator source separation, in our case $1\aas@@fstack{\circ}45$. Rioja et al. (2017) estimate this error to be $\sim 2$ mas. Pradel et al. (2006) obtained a close value of $1.5$ mas via simulated VLBI observations. We used the larger, more conservative value. It should also be noted that the coordinates of the phase-reference calibrator source were measured at $8.4$ GHz, while the EVN observation was conducted at $1.7$ GHz. Frequency-dependent core shift (Sokolovsky et al. 2011) could cause a difference between the International Celestial Reference Frame (ICRF, Charlot et al. 2020) position determined at $8.4$ GHz and the $1.7$-GHz position of the calibrator.
@@
To describe the emission quantitatively, we fitted the visibilities with circular Gaussian brightness distribution model components. Since self-calibration cannot be performed, the obtained flux densities are expected to be lower than the real value due to coherence loss (Martí-Vidal et al. 2010). Previous works (e.g. Mosoni et al. 2006; Gabányi et al. 2019) estimated the coherence loss to be $\sim 25$ %. The parameters of the two fitted components are given in Table 1, where the flux densities have been increased to account for the coherence loss. The errors were calculated according to the formulae given by Fomalont (1999). For the flux density errors, another $10$ % was added in quadrature to account for the absolute amplitude calibration inaccuracies. The full width at half-maximum (FWHM) sizes of the fitted Gaussian features, $\sim 10$ mas and $\sim 4$ mas, corresponding to $\sim 71$ pc and $\sim 28$ pc linear sizes at the source’s redshift, quantify their compact nature.
The brightness temperature of the features can be calculated as (e.g., Veres et al. 2010)
@@
where $S$ is the flux density of the components measured in Jy, $\nu$ the observing frequency in GHz, and $\theta$ the FWHM size in mas. The values obtained are $T_{\mathrm{B}}^{\mathrm{A}}=(4.0\pm 0.8)\cdot 10^{7}$ K and $T_{\mathrm{B}}^{\mathrm{B}}=(8.2\pm 3.2)\cdot 10^{7}$ K for components A and B, respectively. Both values well exceed the upper limit given for star-forming galaxies by Condon (1992), $\sim 10^{5}$ K, indicating AGN-related radio emission.
@@
Figure 4: 5-GHz e-MERLIN image of J1530+1049. The peak intensity is $0.17\mathrm{\,mJy\,beam}^{-1}$. The lowest contours are drawn at $\pm 0.046\mathrm{\,mJy\,beam}^{-1}$ corresponding to $3.5\sigma$ image noise level. Further positive contour levels increase by a factor of $\sqrt{2}$. The restoring beam is shown in the upper left corner, its FWHM size is $121\mathrm{\,mas}\times 36$ mas and its major axis is oriented at a position angle of $23\degr$. The pink crosses and letters indicate the positions of the EVN-detected components A (north), and B (south).
With e-MERLIN, two features (designated as N and S) could be detected at $1.5$ GHz with a separation of $\sim 780$ mas (corresponding to $\sim 5.5$ kpc projected linear size), as shown in Fig. 3. At $5$ GHz, the brighter northern feature (N) is clearly detected with e-MERLIN. At the approximate distance of feature S, at $\sim 700$ mas from N, and in the same position angle as seen in Fig. 3, we recover an emission spot at $\sim 5\sigma$ image noise level at $5$ GHz (Fig. 4).
@@
The results of the model fitting are given in Table 2. The errors have been estimated using the formulae of Fomalont (1999), and additional $20\%$ and $10\%$ were added quadratically to the flux density errors to account for the amplitude calibration uncertainty of e-MERLIN at $1.5$ GHz and $5$ GHz, respectively (Baldi et al. 2021; Horesh et al. 2020).
@@
Figure 5: The flux densities of the two Gaussian components that can be fitted to the e-MERLIN visibilities of J1530++1049. Circles and squares are for the components N and S, respectively. The solid and dashed lines are power-law fits to the corresponding data points. For comparison, the sum of the flux densities of the EVN-detected components is also shown with a yellow upward triangle. Empty downward brown triangles show lower-resolution flux density measurements from various surveys (see Sect. 4.4 for details), and the brown line indicates the power-law fit to those points. The numbers show the corresponding spectral indices.
@@ 4.1 Astrometric registration of the radio components
The same phase-reference source (J1525++1107) was used in all observations. However, it has significant radio emission extended to $\sim 100$-mas scale and the position of its brightness peak can be different at different observing frequencies. To illustrate this, the e-MERLIN images of the calibrator are shown in Fig. 6, with a shift of $\sim 110$ mas between the image cutouts to facilitate their comparison. The north–south oriented double feature seen at $5$ GHz cannot be resolved at $1.5$ GHz with e-MERLIN. However, if the flux densities of the two components relative to each other change with frequency, i.e. the northern feature has higher or very similar flux density than the southern one at the lower frequency of $1.5$ GHz, then the position of the brightness peak can shift towards the northern feature which, on the other hand, is relatively fainter at $5$ GHz. This frequency-dependent change in the location of the brightness peak is further supported by higher-resolution VLBI images of the calibrator taken at $2.3$ GHz, showing an extended bright feature to the north of the mas-scale compact emitting region (the core) at a distance of $\sim(120-150)$ mas. These archival images are available at the Astrogeo website [^4] http://astrogeo.org/cgi-bin/imdb_get_source.csh?source=J1525%2B1107 maintained by L. Petrov. Thus we can conclude that the positional mismatch between the $1.5$-GHz and $5$-GHz e-MERLIN images of J1530++1049 is caused by different brightness distributions of the phase-reference calibrator source at the two frequencies used (Fig. 6). Therefore, in our target source, the brighter northern features detected by e-MERLIN at the two frequencies (Figs. 3 and 4) are co-spatial, and the brighter EVN-detected component (A, Fig. 2) is located within this emitting region. Since the distances and position angles of the target morphology are very similar in the two e-MERLIN images, and no other significant radio emission can be detected in the field at either of the frequencies, we conclude that their cross-identification is solid.